Sheen Co. manufacturers laser printers. It has outlined the following overhead cost drivers:

Overhead Costs Pool Cost Driver Overhead Cost Budgeted Level for
Cost Driver
Quality control Number of inspections $ 73,000 1,000
Machine operation Machine hours 149,000 1,000
Materials handling Number of batches 1,800 40
Miscellaneous overhead cost Direct labor hours 56,000 5,600

Sheen Co. has an order for 1,000 laser printers that has the following production requirements:

Number of inspections 255
Machine hours 220
Number of batches 6
Direct labor hours 730

Using activity-based costing, applied quality control factory overhead for the 1,000 laser printers order is:

Multiple Choice

  • $7,300.

  • $18,155.

  • $18,615.

  • $270.

  • $32,780.
Answer #1

Predetermined overhead rate = Quality control overhead costs / no. Of inspection

= 73000/1000 = $73 per inspection

Applied quality control factory overhead = $73*255 = $18615